What is the pH of a 0.3M acetic acid and 0.3M sodium acetate buffer solution after adding 0.1M NaOH? (Ka = 1.75*10^-5)
Answer #1

Using H-H equation,
pH = pKa + log[sodium acetate]/[acetic acid]
pKa = -logKa = -log(1.75 × 10^-5) = 4.73282827
After adding 0.1 M NaOH,
[sodium acetate] = (0.3 + 0.1) M = 0.4 M
[Acetic acid] = (0.3 - 0.1) M = 0.2 M
pH = 4.73282827 + log (0.4/0.2)
4.73282827 + 0.301029996 = 5.03385827 = 5.03
pH = 5.03